# Explain with Visual Keypoints Like a Real Mentor! A Benchmark for Multimodal Solution Explanation

### v4

@@ Abstract
Figure 1: A student solving a math problem often benefits from visual cues, such as lines, symbols, or highlights, that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions. To serve as effective educational assistants, machines must go beyond answer generation and emulate human-like explanation strategies by explicitly incorporating and referencing visual elements.
@@ 3.1 Benchmark Construction
Average number of VKVK 3.8
@@
Maximum words in $T_{s}^{tldr}$ 123
@@
Maximum words in $T_{s}^{tldr}$ 35.7
@@ 4 Task Definition
Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding. The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem. We compute text similarity using Qwen3 Embedding (qwen3embedding).
@@ 5.1 Toy: Solution Recognition
The multimodal solution explanation tasks are designed to evaluate specific abilities rather than general problem-solving skills. To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal solution explanation task.
@@ Results.
Table 2 shows the accuracy of problem-solving. The 7B generalist baseline struggles, while the 72B model performs second best. Somewhat unexpectedly, the math-specialized models perform worse than the generalist models, likely due to hindered instruction-following capabilities. Among the proprietary baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall. These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal solution explanation task.
@@ 6.1 Qualitative Analysis
Figure 6: Error analysis for Visual Keypoint Identification: Correct choice rates, (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
Visual Keypoint Identification
@@
To gain a finer-grained understanding of the models, we analyze failure cases on a 10% subset of the dataset, as shown in Figure 6. We categorize errors into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements. The three models exhibit similar rates for incorrect descriptions and extra elements but differ substantially in missing elements: Math-PUMA shows the highest rate (38%), followed by Qwen2.5-VL (20%). In contrast, Gemini 2.0 Flash achieves the highest accuracy, while Qwen2.5-VL performs moderately, and Math-PUMA remains the least reliable among the three.
@@
Keypoint-based Explanation Generation
@@ 7 Conclusion
We introduce the multimodal solution explanation task and the ME2 benchmark, which assess multimodal mathematics‐teaching capabilities through two complementary subtasks: identifying essential visual keypoints and generating explanations grounded in those keypoints. Our experimental results demonstrate that current models struggle with both problem solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task. Through our analysis, we find that this limitation stems from the lack of math-specific visual grounding and robust visually grounded reasoning. Enhancing these two abilities will be essential for applying MLLMs effectively in educational settings.
8 Ethics Statement

### v5

@@ Abstract
Figure 1: A student solving a math problem often benefits from visual cues—such as lines, symbols, or highlights—that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions. To serve as effective educational assistants, machines must go beyond answer generation and emulate human-like explanation strategies by explicitly incorporating and referencing visual elements.
@@ 3.1 Benchmark Construction
Average number of VKVK 3.73
@@
Maximum words in $T_{s}^{tldr}$ 198
@@
Maximum words in $T_{s}^{tldr}$ 35.8
@@ 4 Task Definition
Given a problem image $I_{p}$, its text $T_{p}$, the correct answer, and the solution summary $T_{s}^{tldr}$, the model must select, from five candidate sets, the visual keypoints ($VK$) essential for understanding. The four distractor sets are constructed as follows: (1) $VK$ from a problem whose text is semantically similar to $T_{p}$; (2) $VK$ from a problem whose solution summary resembles $T_{s}$; (3) $VK$ from a problem whose own keypoints closely match the target $VK$ (4) $VK$ from a randomly selected problem. Text similarity was computed using Qwen3 Embedding (qwen3embedding), and to ensure reliability, human annotators carefully reviewed and corrected any options exhibiting logical inconsistencies.
@@ 5.1 Toy: Solution Recognition
The multimodal solution explanation tasks are designed to evaluate specific abilities rather than general problem-solving skills. To examine whether the model genuinely understands how to solve problems, we first perform a preliminary study on ME2 prior to the multimodal explanation task.
@@ Results.
Table 2 shows the accuracy of problem-solving. The 7B generalist baseline struggles, while the 72B model performs second best. Somewhat unexpectedly, the math-specialized models perform worse than the generalist models, likely due to hindered instruction-following capabilities. Among the proprietary baselines, GPT-4o struggled similarly to open-source models, whereas Gemini achieved the best performance overall. These results indicate that most MLLMs struggle to recognize the correct solution on ME2, even before performing the multimodal explanation task.
@@ 6.1 Qualitative Analysis
Figure 6: Error analysis for Visual Keypoint Identification: (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
Visual Keypoint Identification.
@@
To gain a finer-grained understanding of the models, we analyze their output behaviors on a 10% subset of the dataset, as shown in Figure 6. We categorize all incorrect outputs into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements. The three models exhibit similar rates for incorrect descriptions and extra elements but differ substantially in missing elements: Math-PUMA shows the highest rate (38%), followed by Qwen2.5-VL (20%). In contrast, Gemini 2.0 Flash achieves the highest accuracy, while Qwen2.5-VL performs moderately, and Math-PUMA remains the least reliable among the three.
@@
Keypoint-based Explanation Generation.
@@ 7 Conclusion
We introduce the multimodal solution explanation task and the ME2 benchmark, which assess multimodal mathematics‐teaching capabilities through two complementary subtasks: identifying essential visual keypoints and generating explanations grounded in those keypoints. Our experimental results demonstrate that current models struggle with both problem-solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task. Through our analysis, we find that this limitation stems from the lack of math-specific visual grounding and robust visually grounded reasoning. Enhancing these two abilities will be essential for applying MLLMs effectively in educational settings.
Ethics Statement